Tickets to the homecoming football game cost $4 for students, $10 for alumni.$12 for other adults, and $6 for children. The total attendance was 3750, and the tickets receipts were $29,100. Six times more students than children attended. The number of alumni was 4/5 of the number of students. How many students, alumni, other adults, and children were at them game?
Answer by richwmiller(17219)

4s+10a+12d+6c=29100,
s+a+d+c=3750,
s=6*c,
a=4s/5
6c+.8(6c)+d+c=3750,
11.8c+d=3750,
4(6c)s+10(.8s)+12d+6c=29100,
78c+12d=29100,
11.8c+d=3750
d=3750-11.8c
78c+12(3750-11.8c)=29100,
63.6c=15900,
c=250
s=1500
a=1200
a+s+c=2950
d=800